OpenStudy (anonymous):

What is the derivative of x^2-2xy+y^3 ?

OpenStudy (unklerhaukus):

Which variable are you taking the derivative with respect to?

OpenStudy (anonymous):

I'm not sure. We're learning Implicit Cost and the first step is to take the derivative, here is the full equation--> x^(2)-2xy+y^(3)=c

OpenStudy (anonymous):

My teacher said to use the product rule and the first step was this 2x-[2y+2xy']+3y^(2)y' but how did he get that?

OpenStudy (unklerhaukus):

Lets take the implicit derivative with respect to x, and see what we get \[f(x)=x^2-2xy+y^3=c\] \[f'(x)=\frac{\mathrm d}{\mathrm dx}(x^2)-2\frac{\mathrm d}{\mathrm dx}(xy)+\frac{\mathrm d}{\mathrm dx}(y^3)=\frac{\mathrm d}{\mathrm dx}(c)%f'(x)=2x-2y-2xy'+3y^2y'\]

OpenStudy (unklerhaukus):

What do you get for this derivative?\[\frac{\mathrm d}{\mathrm dx}(x^2)=\]

OpenStudy (unklerhaukus):

And for these derivatives? \[\frac{\mathrm d}{\mathrm dx}(xy)\] \[\frac{\mathrm d}{\mathrm dx}(y^3)=\] \[\frac{\mathrm d}{\mathrm dx}(c)=\]

OpenStudy (anonymous):

for x^2 the derivative is 2x derivative for xy is y' ? derivative for y^3 is 3y^2 and for c my teacher just plugged in 0. Am I right in the derivatives?

OpenStudy (unklerhaukus):

you have the first term right (x^2)' = 2x

OpenStudy (unklerhaukus):

for the second one you have to use the chain rule \[\frac{\mathrm d}{\mathrm dx}(xy)=\frac{\mathrm d}{\mathrm dx}(x)y+x\frac{\mathrm d}{\mathrm dx}(y)=\]

OpenStudy (unklerhaukus):

\[=\frac{\mathrm dx}{\mathrm dx}y+x\frac{\mathrm dy}{\mathrm dx}=\]

OpenStudy (unklerhaukus):

can you simplify this, using the prime (') notation ?

OpenStudy (unklerhaukus):

(and remember that dx/dx = 1)

OpenStudy (anonymous):

d/dx(xy) x.dy/dx+y

OpenStudy (anonymous):

and derivative of thrid term is 3y^2.dy/dx

OpenStudy (anonymous):

if take derivative of y \[\frac{ d }{ dx }(y)=\frac{ dy }{ dx }\]

OpenStudy (anonymous):

\[f(x)=\frac{ d }{ dx }(x ^{2})-2\frac{ d }{ dx }(xy)+\frac{ d }{ dx }(y ^{3})\]

OpenStudy (anonymous):

Actually I think I'm understanding it a bit more. I just have to stare at it for a while to get it into my head. The derivatives was the only thing that was confusing me. Thank you both for your help :)